OpenStudy (jazzyluvsu123):

is 5/9 + 6/9 = 11/9 ? am I supposed to divide it or anything @SACAPUNTAS @ranga

OpenStudy (anonymous):

Since the denominators the same you can just add straight across. So yes, 5/9 + 6/9 = (5+6)/9 = 11/9

OpenStudy (jazzyluvsu123):

so this is right